# Self-consistent inclusion of disorder in the BCS-BEC crossover near $T_{c}$

### v1

@@ VI Conclusion
In summary, we developed a systematic theoretical approach that incorporates the effects of weak disorder into the BCS-BEC crossover near the critical temperature $T_{c}$. Starting from a functional‐integral formulation in momentum-frequency space, we derived an effective thermodynamic potential by expanding the action to second order in both the disorder potential and the bosonic field, which necessitates inclusion of third‐ and fourth‐order terms in the logarithmic expansion near $T_{c}$. The resulting formalism fully captures Gaussian fluctuations of the order‐parameter field and its coupling to a static white-noise disorder potential. It accurately reproduces the known BCS and BEC limits, recovering the fermionic and bosonic self‐energy diagrams identified in Ref. [19] for the three‐dimensional continuum model, and extends beyond them. The obtained thermodynamic potential and self‐energy expressions remain valid across the entire BCS-BEC crossover, for arbitrarily strong interactions, and consistently reproduce earlier results for non-interacting point bosons and tightly-bound fermion pairs at the lowest order in disorder [27, 19].

### v2

@@ VI Conclusion
In summary, we developed a systematic theoretical approach that incorporates the effects of weak disorder into the BCS-BEC crossover near the critical temperature $T_{c}$. Starting from a functional-integral formulation in momentum-frequency space, we derived an effective thermodynamic potential by expanding the action to second order in both the disorder potential and the bosonic field, which necessitates inclusion of third- and fourth-order terms in the logarithmic expansion near $T_{c}$. The resulting formalism fully captures Gaussian fluctuations of the order-parameter field and its coupling to a static white-noise disorder potential. It accurately reproduces the known BCS and BEC limits, recovering the fermionic and bosonic self-energy diagrams identified in Ref. [20] for the three-dimensional continuum model, and extends beyond them. Within this controlled near-$T_{c}$ framework, the resulting thermodynamic potential and self-energy expressions provide a consistent leading-order description of the BCS-BEC crossover. By recovering established results for non-interacting point bosons and tightly-bound fermion pairs at the lowest order in disorder [30, 20], this approach enables a systematic interpolation between the weak- and strong-coupling regimes. The predicted crossover from disorder-enhanced to disorder-suppressed $T_{c}$ provides a clear benchmark for future cold-atom experiments with tunable speckle disorder. Systematic measurements of the transition temperature across the crossover could therefore directly test the self-consistent fluctuation-disorder interplay developed here.